if 3x + 2y = 19 and 2x + 3y=91, what is the value of x + y?
Answer by checkley77(12844)

3x + 2y = 19 multiply by -2
2x + 3y=91 multiply by 3 & add
-6x-4y=-38
6x+9y=273
-----------------
5y=235
y=235/5
y=47 ans.
3x+2*47=19
3x+94=19
3x=19-94
3x=-75
x=-75/3
x=-25 ans.
Proof:
2*-25+3*47=91
-50+141=91
91=91